what is the pH of the solution that results from 140mL of 0.23 M HF with 230.0 mL of 0.31 M NaF
Answer #1

Solution :-

HF and NaF forms buffer solution

Total volume = 140 ml + 230 ml = 370 ml

initial molarities

HF = 0.23 M    140 ml

NaF = 0.31 M   230 ml

New molarities after mixing

HF = 0.23 M * 140 ml / 370 ml = 0.087 M

NaF = 0.31*230 ml / 370 ml = 0.1927 M

pka of the HF = 3.17

Now using the Henderson equation we can calculate the pH

pH= pka + lof ([base]/ [acid])

pH = 3.17 + log [0.1927 / 0.087]

pH= 3.52

So the pH of the solution is 3.52